A piano tuner stretches a steel piano wire with a tension of 800 N. The steel wire is 0.400 m long and has a mass of 3.00 g. What is the frequency of its fundamental mode of vibration? What is the number of the highest harmonic that could be heard by a person who is capable of hearing frequencies up to 10,000 Hz?


1
Expert's answer
2022-01-21T08:25:52-0500

Explanations & Calculations


  • The harmonics possible for this setup is given by the formula "f_n=n\frac{V}{2L}" .
  • Then the fundamental frequency corresponds to "\small n=1" .
  • The speed of waves on the string is "\small V=\sqrt{\frac{T}{\mu}}"
  • Here "\small \mu": the mass per unit length is "\mu= \frac{0.003\,kg}{0.400\,m}=\small 0.0075\,kgm^{-1}"
  • Therefore,

"\qquad\qquad
\begin{aligned}
\small f_1&=\small 1\times\frac{\sqrt{\frac{800\,N}{0.0075\,kgm^{-1}}}}{2\times0.400\,m}\\
&=\small 408.25\,s^{-1}\,\text{or}\,\,408.25\,Hz
\end{aligned}"


  • This string setup can produce all the consecutive harmonics for "\small n=1,2,3..."
  • Therefore, the highest harmonic would be at some "n" value that would near or equal the maximum hearing frequency.
  • Then,

"\qquad\qquad
\begin{aligned}
\small n\times408.25\,Hz&=\small10000\,Hz\\
\small n&=\small 24.49\\
\small n & \approx \small 24
\end{aligned}"

  • It's the "\small 24^{th}" harmonic.